The point guard of a basketball team has to make a decision about whether or not to shoot a three-point attempt or pass the ball to another player who will shoot a two-point shot. The point guard makes three-point shots 30 percent of the time, while his teammate makes the two-point shot 48 percent of the time.
Xi 3 0
P(xi) 0.30 0.70

Xi 2 0
P(xi) 0.48 0.52



a. What is the expected value for each choice?

b. Should he pass the ball or take the shot himself? Explain.

Answer by stanbon(75887)

The point guard of a basketball team has to make a decision about whether or not to shoot a three-point attempt or pass the ball to another player who will shoot a two-point shot. The point guard makes three-point shots 30 percent of the time, while his teammate makes the two-point shot 48 percent of the time.
Xi 3 0
P(xi) 0.30 0.70
E(x) = 0.3*3 + 0.7*0 = 0.9
---------------------------
Xi 2 0
P(xi) 0.48 0.52
E(x) = 0.48*2 + 0.52*0 = 0.96
----------------------------

a. What is the expected value for each choice?
done
b. Should he pass the ball or take the shot himself? Explain.
Pass the ball ; Expected value is higher.
------------
Cheers,
Stan H.
------------